OpenStudy (anonymous):

1 over 1+tanx + 1 over 1+cotx=1 verify the identities

OpenStudy (goldeverything):

tanx/(1-cotx)= sinx/cosx / (1- cosx/sinx) = sinx/cosx / ((sinx- cosx)/sinx) =sin^2x/ cosx( sinx-cosx) Similarly cotx/(1-tanx)= cos^x/sinx(cosx-sinx) LHS=(tan x / (1 - cot x)) + (cot x / (1 - tan x)) = sin^2x/ cosx( sinx-cosx)+cos^x/sinx(cosx-sinx) =(sin^3x-cos^x)/(sin x cosx(sinx-cosx)) =(sin^2x+sinx cosx+cos^2x)/ sinx cosx =( 1+ sinx cosx)/ sinx cosx = 1/ sinx cosx +1 =1+ 1/sinx cosx =1+ sec x cosex